OpenStudy (lisa123):

solve for x. 4x^2=-13x-3

OpenStudy (campbell_st):

rewrite the problem as \[4x^2 + 13x + 3 = 0\] ' so to solve this I multiply 4 and 3 then find the factors that add to 13 any ideas..?

OpenStudy (lisa123):

12 and 1

OpenStudy (campbell_st):

great so next write the factorisation as \[\frac{(4x + 12)(4x + 1)}{4}\] then remove the common factors from the numerators \[\frac{4(x + 3)(4x + 1)}{4} = 0\] cancel out the common factor in the numerator ad denominator then you can find the solutions

OpenStudy (lisa123):

so (x+3)(4x+1)

OpenStudy (campbell_st):

that's the factored for so you have x +3 = 0 or 4x + 1 = 0 so find the 2 values of x that make the equations true

OpenStudy (lisa123):

-3 and -1/4

OpenStudy (campbell_st):

that's correct